Just wondering if someone could explain how to solve to find x:
2 × (0.1)^x+1 = 25
thanking you
Tracey

Answer by lwsshak3(11628)

ust wondering if someone could explain how to solve to find x:
2 × (0.1)^x+1 = 25
**
2 × (0.1)^x+1 = 25
2*(0.1)^x=25-1=24
(0.1)^x=24/2=12
take log of both sides
xlog(0.1)=log12
x=log12/log(0.1)
x=1.0792/-1
x=-1.0792